OpenStudy (anonymous):

Find all vertical asymptotes, x=a, of the following functions. For each value of a, evaluate lim x->a^+,lim x->a^-,and limx->a

OpenStudy (anonymous):

\[f(x) = \frac{ x^3-10x^2+16x }{ x^2-8x }\]

OpenStudy (anonymous):

(a) Find all vertical asymptotes x=a (b)Evaluate limits at \[\lim_{x \rightarrow a^+}\] \[\lim_{x \rightarrow a^-}\] \[\lim_{x \rightarrow a}\]

OpenStudy (anonymous):

I know that to find a vertical asymptote, \[h(x)=\frac{ f(x) }{ g(x) }\] where if the demnominatopr at g(x) is 0 and the numerator at f(c) is not zero than h is a vertical asymptote where x=c.

OpenStudy (anonymous):

Factoring out f(x): \[f(x) = \frac{x^3-10x^2+16x} {x^2-8x}\] \[=\frac{x(x-2)(x-8)}{x(x-8)}\] When x = 8 g(x) is zero and f(x) is not zero so I would have a vertical asymptote at x=8

OpenStudy (anonymous):

Is this right so far?